Calculate how much 10% salt solution and 50% salt solution must be mixed to end up with exactly 20 gallons of a 40% salt solution.
Answer by stanbon(75887)

Calculate how much 10% salt solution and 50% salt solution must be mixed to end up with exactly 20 gallons of a 40% salt solution.
---------
Equation:
salt + salt = salt
0.10x + 0.50(20-x) = 0.40*20
------
Multiply thru by 100 to get:
10x + 50*20 - 50x = 40*20
-40x = -10*20
x = 5 gallons (amt. of 10% solution needed)
-----
20-x = 15 gallons (amt. of 50% solution needed)
============================
Cheers,
Stan H.
===========